# The Dirichlet problem for hessian quotient type curvature equations in Minkowski space

### v1

@@ 4. $C^{2}$ estimates
Suppose bounded domain $\Omega\subset\mathbb{R}^{n}$ is $(\eta,k)$-convex with smooth boundary $\partial\Omega$. Let $u\in C^{3}(\overline{\Omega})$ be an admissible solution of (1.3). Then there exists a positive constant $C$ depending only on $\theta_{0}$, $\|\psi\|_{C^{1}(\overline{\Omega}\times[-\mu_{0},\mu_{0}])}$, $\|\varphi\|_{C^{3}(\overline{\Omega})}$ and $\partial\Omega$ satisfying
@@
(4.2) $x_{n}=\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha<n}\kappa^{b}_{\alpha}x_{\alpha}^{2}+O(|x^{\prime}|^{3}),$
where $\kappa^{b}_{1},\ldots,\kappa^{b}_{n-1}$ are the principal curvatures of $\partial\Omega$ at the origin and $x^{\prime}=(x_{1},\ldots,x_{n-1})$. Define
@@
(4.3) $|u_{\alpha\beta}(0)|\leqslant C\mbox{ for }1\leqslant\alpha,\beta\leqslant n-1,$
@@
Since $\Omega$ is $(\eta,k)$-convex, there exist two positive constants $\theta$ and $K$ satisfying
(4.9) $(\kappa_{1}^{b}-3\theta,\ldots,\kappa_{n-1}^{b}-3\theta,2K)\in\widetilde{\Gamma}_{k}.$
@@
on $\overline{\omega}_{\delta}$. We see that when $\delta$ depending on $\theta$ and $K$ is sufficiently small, we have
[…]
In view of (4.2) and (4.9), $\lambda(D^{2}v)\in\widetilde{\Gamma}_{k}$ on $\overline{\omega}_{\delta}$. Thus, there exists an uniform constant $\eta_{0}>0$ depending only on $\theta$, $\partial\Omega$ and $K$ satisfying
@@
Actually, we can assume $D^{2}v=\operatorname{diag}(\kappa_{1}^{b}-2\theta,\dots,\kappa_{n-1}^{b}-2\theta,2K)$ by rotating the coordinate system if necessary and note that $\gamma^{ij}=\operatorname{diag}(1,\dots,1,\frac{1}{w})$ at 00 since $|Du(0)|=u_{n}(0)$. It follows that
$\{\frac{1}{w}(\gamma^{ik}D_{kl}v\gamma^{lj})\}=\operatorname{diag}(\frac{\kappa_{1}^{b}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{b}-2\theta}{w},\frac{2K}{w^{3}}).$
Let $\kappa_{v}=(\frac{\kappa_{1}^{b}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{b}-2\theta}{w},\frac{2K}{w^{3}})=(\kappa_{v}^{\prime},\frac{2K}{w^{3}})$, we have
$\frac{\sigma_{k}}{\sigma_{l}}(\kappa_{v})>\frac{\sigma_{k}}{\sigma_{l}}(\kappa_{v}^{\prime},\frac{2K}{w})>0$
by $w<1$, (4.9) and (2.6). Then we can find a suitable constant $\delta$ such that (4.12) holds.
@@
where $v(x)$ is defined by (4.10), $d(x):=\mathrm{dist}(x,\partial\Omega)$ is the distance from $x$ to the boundary $\partial\Omega$, and $t,N$ are two positive constants to be determined later. Since $f^{1/(k-l)}$ is concave in $\Gamma$ and homogeneous of degree one and $|Dd|\equiv 1$ on the boundary $\partial\Omega$, by (4.12) and (2.3), we have,
@@
(4.17) Φ0 on $\Phi\leqslant 0\mbox{ on }\overline{\omega}_{\delta}$
by choosing suitable positive constants $\delta$, $t$, $N$ and $R$.
@@
(4.19) $-td+\frac{N}{2}d^{2}\leqslant 0\mbox{ on }\overline{\omega}_{\delta}.$
@@
$a_{ij}=\begin{pmatrix}\frac{u_{11}}{\omega}&0&\cdots&0&\frac{u_{1n}}{\omega}\\ 0&\frac{u_{22}}{\omega}&\cdots&0&\frac{u_{2n}}{\omega}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\frac{u_{n-1,n-1}}{\omega}&\frac{u_{n-1,n}}{\omega}\\ \frac{u_{n1}}{\omega}&\frac{u_{n2}}{\omega}&\cdots&\frac{u_{n,n-1}}{\omega}&\frac{u_{nn}}{\omega^{3}}\end{pmatrix}$
@@
$\begin{cases}\kappa_{\alpha}=\dfrac{u_{\alpha\alpha}}{\omega}+o(1),&\alpha=1,\cdots,n-1,\\[10.0pt] \kappa_{n}=\dfrac{u_{nn}}{\omega^{3}}+O(1),\end{cases}$
@@
Since k ¡ n, by choosing $R_{1}$ large enough, we have at the origin,
$\psi=\Big{(}\frac{\sigma_{k}}{\sigma_{l}}\Big{)}(\lambda(\eta))=\frac{\big{(}C_{n}^{k}-(n-k+1)\big{)}u_{nn}^{k}/(w^{3k})+O(u_{nn}^{k-1})}{(C_{n}^{l}-(n-l+1))u_{nn}^{l}/(w^{3l})+O(u_{nn}^{l-1})}.$
@@
$-G^{s}W_{s}\leqslant C|DW|+\frac{\epsilon_{0}}{2}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{0}}G^{ij}W_{i}W_{j}.$
@@
$=\omega\kappa_{r}(\eta_{\alpha r}+\rho_{\alpha}\eta_{nr}-\sum_{\beta\leqslant n-1}u_{\beta}\eta_{\beta r})+\gamma^{sj}b_{jr}\rho_{\alpha s}u_{n}.$
@@
(4.35) $|\gamma^{sj}b_{jr}W_{s}|\leqslant C\omega\kappa_{r}(\epsilon+|\rho_{\alpha}|)+C.$
@@
$\frac{2}{\omega}f_{r}\kappa_{r}|(\sum_{t}b_{tr}u_{t})\gamma^{sj}b_{jr}W_{s}|=$ $\frac{2}{\omega}((k-l)\psi-\sum_{i\neq r}f_{i}\kappa_{i})\left|\left(\sum_{t}b_{tr}u_{t}\right)\gamma^{sj}b_{jr}W_{s}\right|$
@@
$\leqslant$ $C\psi|DW|+(C\epsilon_{0}^{-1}(\epsilon+|\rho_{\alpha}|)+\epsilon_{1})\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{1}}\sum_{i}G^{ii}$
@@
where $m=1-\theta^{2}$, $\theta$ is defined in spacelike. Therefore, as in Case (b-1), (4.8) follows by (4.27). ∎
@@
(4.36) $F^{ij}\nabla_{i}\nabla_{j}h_{ab}=$ $-F^{ij,kl}\nabla_{a}h_{ij}\nabla_{b}h_{kl}-F^{ij}h_{ij}h_{ak}h_{bk}$
$+F^{ij}h_{ik}h_{jk}h_{ab}+\nabla_{a}\nabla_{b}\psi$
@@
(4.37) $0\geqslant\tilde{F}^{ii}\nabla_{ii}h_{11}=-\tilde{F}^{ij,kl}\nabla_{1}h_{ij}\nabla_{1}h_{kl}-\tilde{\psi}h_{11}^{2}+h_{11}\tilde{F}^{ii}h_{ii}^{2}+\nabla_{11}(\tilde{\psi}).$
@@
Proof of theorem 1.3.
Combined (3.2), (3.1), (4.38), we can obtain
(4.39) $||u||_{C^{2}(\overline{\Omega})}\leqslant C.$
Applying the Evans-Krylov theory, we can obtain the $C^{2,\alpha}$ estimates for $u$. Using the continuity method and the maximum principle, we can finally establish the existence and uniqueness of solutions to (1.3). ∎

### v2

@@ 4. $C^{2}$ estimates
Suppose $\Omega\subset\mathbb{R}^{n}$ is a bounded domain with smooth boundary $\partial\Omega$. Let $u\in C^{3}(\overline{\Omega})$ be an admissible solution of (1.3). Then there exists a positive constant $C$ depending only on $\theta_{0}$, $\|\psi\|_{C^{1}(\overline{\Omega}\times[-\mu_{0},\mu_{0}])}$, $\|u\|_{C^{1}(\overline{\Omega})}$ and $\partial\Omega$ satisfying
@@
(4.2) $x_{n}=\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha<n}\kappa^{\prime}_{\alpha}x_{\alpha}^{2}+O(|x^{\prime}|^{3}),$
where $\kappa^{\prime}_{1},\ldots,\kappa^{\prime}_{n-1}$ are the principal curvatures of $\partial\Omega$ at the origin as before and $x^{\prime}=(x_{1},\ldots,x_{n-1})$. Define
@@
(4.3) $|u_{\alpha\beta}(0)|\leqslant C\mbox{ for }1\leqslant\alpha,\beta\leqslant n-1.$
@@
Note that there exist positive constants $\theta$ and $K$ such that
(4.9) $(\kappa_{1}^{\prime}-3\theta,\ldots,\kappa_{n-1}^{\prime}-3\theta,2K)\in\widetilde{\Gamma}_{k}$
holds.
@@
on $\overline{\omega}_{\delta}$. We see that when $\delta(\theta,K)$ is sufficiently small, we have
(4.11) $v\leqslant$ $-\frac{\theta}{2}|x^{\prime}|^{2},$ $\mbox{ on }\partial_{1}\omega_{\delta},$
$v\leqslant$ $-\frac{\delta^{2}}{2},$ $\mbox{ on }\partial_{2}\omega_{\delta},$
$v\leqslant$ $-\frac{\theta\delta^{2}}{2},$ $\mbox{ on }\partial_{3}\omega_{\delta}.$
Besides, by (4.2) and (4.9), we have
(4.12) $\lambda(D^{2}v)=(\kappa_{1}^{\prime}-2\theta,\ldots,\kappa_{n-1}^{\prime}-2\theta,2K)\in\widetilde{\Gamma}_{k}$
on $\overline{\omega}_{\delta}$, where $\lambda(D^{2}v)$ denotes the eigenvalues of $D^{2}v$. Thus, there exists an uniform constant $\eta_{0}>0$ depending only on $\theta$, $\partial\Omega$ and $K$ satisfying
@@
Actually, we can assume $D^{2}v=\operatorname{diag}(\kappa_{1}^{\prime}-2\theta,\dots,\kappa_{n-1}^{\prime}-2\theta,2K)$ by rotating the coordinate system if necessary and note that $\gamma^{ij}=\operatorname{diag}(1,\dots,1,\frac{1}{w})$ at 0 since $|Du(0)|=u_{n}(0)$. It follows that
$\left\{\frac{1}{w}(\gamma^{ik}D_{kl}v\gamma^{lj})\right\}=\operatorname{diag}\left(\frac{\kappa_{1}^{\prime}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{\prime}-2\theta}{w},\frac{2K}{w^{3}}\right).$
Let $\kappa_{v}=(\frac{\kappa_{1}^{\prime}-2\theta}{w},\dots,\frac{\kappa_{n-1}^{\prime}-2\theta}{w},\frac{2K}{w^{3}})=(\kappa_{v}^{\prime},\frac{2K}{w^{3}})$. By (4.12) and $w<1$, we have $(\kappa_{v}^{\prime},\frac{2K}{w})\in\widetilde{\Gamma}_{k}$. Then it follows from the ellipticity of $\sigma_{k}$ in $\widetilde{\Gamma}_{k}$ that $\kappa_{v}\in\widetilde{\Gamma}_{k}$. Hence we can find a suitable constant $\delta$ such that (4.13) holds.
@@
where $v(x)$ is defined by (4.10), $d(x):=\mathrm{dist}(x,\partial\Omega)$ is the distance from $x$ to the boundary $\partial\Omega$, and $t,N$ are two positive constants to be determined later. Since $f^{1/(k-l)}$ is concave in $\widetilde{\Gamma}_{k}$ and homogeneous of degree one and $|Dd|\equiv 1$ on the boundary $\partial\Omega$, by (4.13) and (2.4), we have,
@@
(4.18) $\Phi\leqslant 0\mbox{ on }\overline{\omega}_{\delta}$
by choosing suitable positive constants $\delta$, $t$, NN and $R$.
@@
(4.20) $td+\frac{N}{2}d^{2}\leqslant 0\mbox{ on }\overline{\omega}_{\delta}.$
@@
$a_{ij}=\begin{pmatrix}\frac{u_{11}}{w}&0&\cdots&0&\frac{u_{1n}}{w}\\ 0&\frac{u_{22}}{w}&\cdots&0&\frac{u_{2n}}{w}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\frac{u_{n-1,n-1}}{w}&\frac{u_{n-1,n}}{w}\\ \frac{u_{n1}}{w}&\frac{u_{n2}}{w}&\cdots&\frac{u_{n,n-1}}{w}&\frac{u_{nn}}{w^{3}}\end{pmatrix}$
@@
$\begin{cases}\kappa_{\alpha}=\dfrac{u_{\alpha\alpha}}{w}+o(1),&\alpha=1,\cdots,n-1,\\[10.0pt] \kappa_{n}=\dfrac{u_{nn}}{w^{3}}+O(1),\end{cases}$
@@
Since $k<n$, by choosing $R_{1}$ large enough, we have at the origin,
$\psi=\Big(\frac{\sigma_{k}}{\sigma_{l}}\Big)(\lambda(\eta))=\frac{C_{n-1}^{k}u_{nn}^{k}/(w^{3k})+O(u_{nn}^{k-1})}{C_{n-1}^{l}u_{nn}^{l}/(w^{3l})+O(u_{nn}^{l-1})}.$
@@
$G^{s}W_{s}\leqslant C|DW|+\frac{\epsilon_{0}}{2}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{0}}G^{ij}W_{i}W_{j}.$
@@
$=w\kappa_{r}(\eta_{\alpha r}+\rho_{\alpha}\eta_{nr}-\sum_{\beta\leqslant n-1}u_{\beta}\eta_{\beta r})+\gamma^{sj}b_{jr}\rho_{\alpha s}u_{n}.$
@@
(4.36) $|\gamma^{sj}b_{jr}W_{s}|\leqslant Cw\kappa_{r}(\epsilon+|\rho_{\alpha}|)+C.$
@@
$\frac{2}{w}f_{r}\kappa_{r}|(\sum_{t}b_{tr}u_{t})\gamma^{sj}b_{jr}W_{s}|=$ $\frac{2}{w}((k-l)\psi-\sum_{i\neq r}f_{i}\kappa_{i})\left|\left(\sum_{t}b_{tr}u_{t}\right)\gamma^{sj}b_{jr}W_{s}\right|$
@@
$\leqslant$ $C|DW|+(C\epsilon_{0}^{-1}(\epsilon+|\rho_{\alpha}|)+\epsilon_{1})\sum_{i\neq r}f_{i}\kappa_{i}^{2}+\frac{C}{\epsilon_{1}}\sum_{i}G^{ii}$
@@
where $m=1-\theta_{0}^{2}$, $\theta_{0}$ is defined in spacelike. Therefore, as in Case (b-1), (4.8) follows by (4.28). ∎
@@
(4.37) $F^{ij}\nabla_{ij}h_{ab}=$ $-F^{ij,kl}\nabla_{a}h_{ij}\nabla_{b}h_{kl}-F^{ij}h_{ij}h_{ak}h_{bk}$
$+F^{ij}h_{ik}h_{jk}h_{ab}+\nabla_{ab}\psi$
@@
[…]
@@
We have done.
5. Proof of the main result
The proof of Theorem 1.3 relies on a standard continuity method [5]. Let
$\psi^{0}(x,\underline{u})=\frac{\sigma_{k}}{\sigma_{l}}\left(\lambda(\eta[M_{\underline{u}}])\right)$
and extend $\psi^{0}$ to $\Omega\times\mathbb{R}$ with $\psi^{0}(x,t)=\psi^{0}(x,\underline{u})$. Extend $\psi$ in the same manner, we consider, for each $t\in[0,1]$,
(5.1) $\left\{\begin{aligned} &\frac{\sigma_{k}}{\sigma_{l}}\left(\lambda(\eta[M_{u^{t}}])\right)=t\psi+(1-t)\psi^{0}&&\mbox{in}~\Omega,\\ &u^{t}=0&&\mbox{on}~\partial\Omega.\end{aligned}\right$.
The set $S$ consists of numbers $t\in[0,1]$ such that there exists a function $u^{t}\in C^{2,\alpha}(\overline{\Omega})$, which is an admissible solution of (5.1).
Obviously, $\underline{u}$ is the solution of (5.1) for $t=0$, so the set $S$ is nonempty. We shall prove that $S$ is both open and close, so $S=[0,1]$ and the proof is finished. First, by the implict function theorem, we can obtain $S$ is an open set. Next, let $\{s_{i}\}_{i\in\mathbb{N}}$ be an arbitrary sequence from the set $S$ and $s\in[0,1]$ satisfies $\lim_{i\to\infty}s_{i}=s$, we shall prove $s\in S$. Combine with the estimates
$\sup_{\overline{\Omega}}|Du^{t}|\leqslant 1-\theta_{0}<1$
and
$\|u^{t}\|_{C^{2}(\overline{\Omega})}\leqslant C,$
we obtain the uniform ellipticity of $\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta))$.
Since $(\frac{\sigma_{k}}{\sigma_{l}}(\lambda(\eta)))^{\frac{1}{k-l}}$ is concave (see [28]) in $\widetilde{\Gamma}_{k}$, then by Evans–Krylov theorem, we can obtain a $C^{2,\alpha}$ estimate, which means
$\|u^{s_{i}}\|_{C^{2,\alpha}(\overline{\Omega})}\leqslant C,$
where $C$ is independent on $s_{i}$. By the Arzela-Ascoli theorem, we have $u^{s}\in C^{2,\delta}(\overline{\Omega})$ for $\delta\in(0,\alpha)$. Therefore, $s\in S$ and $S$ is closed.
In addition, by standard nonlinear elliptic regularity, we can obtain $u\in C^{\infty}(\overline{\Omega})$.
The uniqueness of the solution follows from the maximum principle.
Acknowledgment. We would like to thank Professor Heming Jiao and the referees for helpful comments and valuable suggestions.